The heights of 9-year-old children are normally distributed, with a mean of 123 cm and a standard deviation of 10 cm. Find the probability that the mean height of three randomly selected 9-year-old children is greater than 125 cm


1
Expert's answer
2021-02-04T08:09:58-0500

"P(X>125)"


for "X=125, Z=(125-123)\/10"


Z=0.2


"\therefore P(X>125)=P(Z>0.2)"


"=1-0.5080"


"=0.492"